In 13 years i will be three times the age as i was 13 years ago. What will be my age ?
Answer by Edwin McCravy(20059)

A = present age
A+13 = age in 13 years
A-13 = age 13 years ago

A+13 = 3(A - 13)

You finish.

Edwin